Determine if
f(x)=ln(5x+4), x>−4/5, has an inverse, and if it does, find this inverse.

Answer by lwsshak3(11628)

Determine if
f(x)=ln(5x+4), x>−4/5, has an inverse, and if it does, find this inverse
**
y=ln(5x+4)
Change to exponential form: base(e) raised to log of number(y)=number(5x+4)
e^y=5x+4
interchange x and y, then solve for y, the inverse.
e^x=5y+4
5y=e^x-4
y=(e^x-4)/5
f(x) inverse=(e^x-4)/5